The following data is a part of the California district test score average and student-teacher-ratio.

Observation Number

TS

STR

1

690.8

17.89

2

661.2

21.52

3

643.6

18.70

4

647.7

17.36

5

640.8

18.67

6

605.6

21.41

7

606.8

19.50

8

609.0

20.89

9

612.5

19.95

10

612.7

20.81

11

615.8

21.24

12

616.3

21.00

13

616.3

20.60

14

616.3

20.01

15

616.5

18.03

16

617.3

20.25

17

618.1

16.98

18

618.3

16.51

19

619.8

22.70

20

620.3

19.91

Using the data given we want to retrieve a linear regression line that reflects the effect of STR on TS

  1. Calculate OLS estimator of TS.
  2. Construct 95% confidence interval for Test score and Student teacher ratio
Answer #1

a)

Observation Number TS (y) STR (x) xi^2 yi^2 xi*yi
1 690.8 17.89 320.0521 477204.6 12358.41
2 661.2 21.52 463.1104 437185.4 14229.02
3 643.6 18.7 349.69 414221 12035.32
4 647.7 17.36 301.3696 419515.3 11244.07
5 640.8 18.67 348.5689 410624.6 11963.74
6 605.6 21.41 458.3881 366751.4 12965.9
7 606.8 19.5 380.25 368206.2 11832.6
8 609 20.89 436.3921 370881 12722.01
9 612.5 19.95 398.0025 375156.3 12219.38
10 612.7 20.81 433.0561 375401.3 12750.29
11 615.8 21.24 451.1376 379209.6 13079.59
12 616.3 21 441 379825.7 12942.3
13 616.3 20.6 424.36 379825.7 12695.78
14 616.3 20.01 400.4001 379825.7 12332.16
15 616.5 18.03 325.0809 380072.3 11115.5
16 617.3 20.25 410.0625 381059.3 12500.33
17 618.1 16.98 288.3204 382047.6 10495.34
18 618.3 16.51 272.5801 382294.9 10208.13
19 619.8 22.7 515.29 384152 14069.46
20 620.3 19.91 396.4081 384772.1 12350.17
Total 12505.7 393.93 7813.52 7828232 246109.5
Average 625.285 19.6965

Sxx 54.47725
Syy 8605.366
Sxy -209.029

let the ols b y = b0 + b1x

b1 = Sxy/Sxx = -209.029/54.477 = -3.837

b0 = = 625.285 - (-3.837*19.697) = 700.86

y = 700.86 - 3.837*x

The ordinary least squares regression line will be

TS = 700.86 - 3.837*STR

b) 95% CI calculated from minitab output

One-Sample T: TS (y), STR (x)

Descriptive Statistics

Sample N Mean StDev SE Mean 95% CI for μ
TS (y) 20 625.28 21.28 4.76 (615.32, 635.25)
STR (x) 20 19.697 1.693 0.379 (18.904, 20.489)